Grade 12th passMechanics

the motion of a body falling from rest in a medium is given by equation, dv/dt = 1- 2v. the velocity at any time t is

nandini
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago
dv)/dt = 1-2v

dv (1/(1-2v) = dt

Now,

integrating both sides we get,

(-1/2)×ln(1- 2v) = t .....Zero initial conditions ,hence c=0

ln(1- 2v) = -2t

1 -2v = e^-2t

v = (1 - e^-2t)/2

This is the required velocity time relation